# AtlasD: Automatic Local Symmetry Discovery

### v1

@@ 6 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgement
Appendix A Proofs
A.1 Atlas equivariance of gauge equivariant CNN
@@
We first clarify a few definitions. By trivial charts, we simply mean that $\varphi_{c}$ is the inclusion map. We model a gauge equivariant CNN as a series of convolutional layers and pointwise nonlinearities (Cohen et al., 2019). To enforce gauge equivariance, we will require all kernels $K:\mathbb{R}^{\text{d}}\to\mathbb{R}^{C_{\text{in}}\times C_{\text{out}}}$ in the network to satisfy $K(g^{-1}v)=K(v)$ for $g\in G$.
@@
A.2 Argmin of Standard Basis Regularization
@@
Suppose $\{b_{i}\}$ is a basis of $V$. If $b$ is disjoint then for all $i\neq j$ we have $|b_{i}|\cdot|b_{j}|=0$ since for each index $\ell$, at least one of $b_{i,\ell}$ or $b_{j,\ell}$ will be zero. Conversely, if $b$ is not disjoint, then there exist some $i\neq j$ such that $|b_{i}|\cdot|b_{j}|>0$. To see this, note that we must have some $i\neq j$ where $b_{i}$ and $b_{j}$ share a non-zero element at index $\ell$. $|b_{i}|\cdot|b_{j}|$, the sum of non-negative numbers, is then greater or equal to $|b_{i,\ell}||b_{j,\ell}|>0$.
@@ C.1 Global Symmetry Discovery
For coset discovery, we seed the model with $K=256$ basis elements and run $3$ epochs with cross-entropy loss. We believe $K$ can be reduced significantly and still discover the parity component. However, apart from computational time, we do not observe any downsides to keeping the number high. To filter out the final representatives, we find all the unique cosets in the top $q=16$ matrices. The learning rate is $0.001$.
@@ Appendix D Additional Experiments
D.1 Coset Discovery on a Synthetic Problem
We want to see if ATLAS can discover multiple cosets in situations where the loss of a one coset (typically the identity) is lower than the loss of another coset in the symmetry group. This could happen when the trained predictor is not super accurate or if the symmetry of the true function itself is not perfect. To do so, we experiment on the global discrete symmetries of the function $f(x,y)=\arctan\left(\frac{y+0.1}{x}\right)$. The identity transformation is a perfect symmetry of $f$ and rotation by $180$ degrees is an approximate symmetry. Although there is some variability between runs, Figure 11 highlights we are able to consistently discover rotation coset representatives among the top $24$ cosets.
@@ D.4 Comparison with LieGG in PDE Experiment
Figure 15: Singular values of generators discovered by LieGG in modifed PDE experiment. We run twice: once without the rectangular heat source (blue) and once with the heat source (red).
@@
For simplicity, we consider the setting where LieGG is given access to the ground truth partial differential equation instead of a predictor network. We also do not perform time stepping and focus solely on the global symmetries of $\frac{\partial u}{\partial t}=\alpha(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}})$. This setup, while slightly different from the experiments for our method, only makes it easier for LieGG to learn the symmetry. The singular values of the discovered generators are shown in Figure 15 .

### v2

@@ 6 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Acknowledgement
Appendix A Math
A.1 Connection between Sheaf Morphism and Atlas Locality
Sheaves provide a framework for relating local data across overlapping regions and are relevant in the context of local functions. In particular, we show that being a sheaf morphism is a strictly stronger condtion than being atlas local. We first summarize the definitions related to sheaves and sheaf morphisms, and refer to Vakil (2022) for a more detailed introduction.
Suppose that $X$ is a topological space. $\mathcal{F}$ is said to be a presheaf (of sets) on $X$ if the following conditions are met:
  1. 1.
    For any open set $U\subseteq X$, we have a set $\mathcal{F}(U)$ containing the sections of $F$ over $U$.
  2. 2.
    If $U$ and $V$ are open sets such that $U\subseteq V$, we have a restriction function $\mathrm{res}_{V,U}:\mathcal{F}(V)\to\mathcal{F}(U)$.
  3. 3.
    $\mathrm{res}_{U,U}$ must be the identity.
  4. 4.
    For any triplet of open sets such that $U\subseteq V\subseteq W$, we have $\mathrm{res}_{W,U}=\mathrm{res}_{V,U}\circ\mathrm{res}_{W,V}$.
Then, any presheaf $\mathcal{F}$ is said to be a sheaf (of sets) if the additional two conditions are satisfied:
  1. 1.
    Let $I$ be an indexing set and suppose we have an open cover $\{U_{i}\}_{i\in I}$ of $U$ where $U_{i}\subseteq U$. For any $f_{1},f_{2}\in\mathcal{F}(U)$, if $\mathrm{res}_{U,U_{i}}\ f_{1}=\mathrm{res}_{U,U_{i}}\ f_{2}$ for all $i\in I$, we must have $f_{1}=f_{2}$
  2. 2.
    For any open cover $\{U_{i}\}_{i\in I}$ of $U$ and family of sections $\{f_{i}\in U_{i}~{}|~{}i\in I\}$, if $\mathrm{res}_{U_{i},U_{i}\cap U_{j}}\ f_{i}=\mathrm{res}_{U_{j},U_{i}\cap U_{j}}\ f_{j}$ for all $i,j\in I$ , then there must exist $f\in\mathcal{F}(U)$ so that $\mathrm{res}_{U,U_{i}}\ f=f_{i}$ for all $i$.
Finally, a sheaf morphism between the sheaves $\mathcal{F}$ and $\mathcal{G}$ is a morphism $\phi:\mathcal{F}\to\mathcal{G}$ specified by a map $\phi_{U}:\mathcal{F}(U)\to\mathcal{G}(U)$ for each open set $U\subseteq X$ so that $\phi$ commutes with restriction, i.e., if $V\subseteq U$ we have $\mathrm{res}^{\mathcal{G}}_{U,V}\ \phi_{U}(\mathcal{F}(U))=\phi_{V}(\mathrm{res}^{\mathcal{F}}_{U,V}\ \mathcal{F}(U))$
Now, suppose we have a manifold ${\mathcal{M}}$ and an atlas specified by a collection of charts $\{U_{c}\}_{c=1}^{n}$. We may consider ${\mathcal{M}}$ as a topological space whose subbasis is given by the collection of charts. Note that for any fixed $d$, the set of feature fields specified by functions $F:{\mathcal{M}}\to\mathbb{R}^{d}$ forms a sheaf. In particular, let $\phi$ be a sheaf morphism between feature fields of dimension $d_{\text{in}}$ to those of dimension $d_{\text{out}}$. Then, for each $U_{c}$ and arbitrary $F:{\mathcal{M}}\to\mathbb{R}^{d_{\text{in}}}$, we have that $\mathrm{res}^{\text{out}}_{{\mathcal{M}},U_{c}}\ \phi_{\mathcal{M}}(F)=\phi_{U_{c}}(\mathrm{res}^{\text{in}}_{{\mathcal{M}},U_{c}}\ F)$ for all $U_{c}$. To finally prove $\Phi=\phi_{\mathcal{M}}$ is atlas local, we can construct each $\Phi_{c}$ using $\phi_{U_{c}}$. This is because for any chart $U_{c}$, the sheaf morphism condition guarantees that the local behavior of $\phi_{\mathcal{M}}$ is fully captured by $\phi_{U_{c}}$.
However, unlike sheaf morphisms, atlas locality does not require a localized map on all open subsets, e.g., the intersection of charts, and is thus a weaker condition.
A.2 Atlas equivariance of gauge equivariant CNN
@@
We first clarify a few definitions. By trivial charts, we mean that $\varphi_{c}$ is the inclusion map. We model a gauge equivariant CNN as a series of convolutional layers and pointwise nonlinearities (Cohen et al., 2019). To enforce gauge equivariance, we will require all kernels $K:\mathbb{R}^{\text{d}}\to\mathbb{R}^{C_{\text{in}}\times C_{\text{out}}}$ in the network to satisfy $K(g^{-1}v)=K(v)$ for $g\in G$.
@@
A.3 Argmin of Standard Basis Regularization
@@
Suppose $\{b_{i}\}$ is a basis of $V$. If $b$ is disjoint then for all $i\neq j$ we have $|b_{i}|\cdot|b_{j}|=0$ since for each index $\ell$, at least one of $b_{i,\ell}$ or $b_{j,\ell}$ will be zero. Conversely, if $b$ is not disjoint, then there exist some $i\neq j$ such that $|b_{i}|\cdot|b_{j}|>0$. To see this, note that we must have some $i\neq j$ where $b_{i}$ and $b_{j}$ share a non-zero element at index $\ell$. $|b_{i}|\cdot|b_{j}|$, the sum of non-negative numbers, is then greater than or equal to $|b_{i,\ell}||b_{j,\ell}|>0$.
@@ C.1 Global Symmetry Discovery
For coset discovery, we seed the model with $K=64$ basis elements and run $3$ epochs with cross-entropy loss. To filter out the final representatives, we find all the unique cosets in the top $q=16$ matrices. The learning rate is $0.001$.
@@ Appendix D Additional Experiments
D.1 Coset Discovery on Imperfect Data
We want to see if AtlasD can discover multiple cosets in situations where the loss of a one coset (typically the identity) is lower than the loss of another coset in the symmetry group. This could happen when the trained predictor is not super accurate or if the symmetry of the true function itself is not perfect. To do so, we experiment on the global discrete symmetries of the function $f(x,y)=\arctan\left(\frac{y+0.1}{x}\right)$. The identity transformation is a perfect symmetry of $f$ and rotation by $180$ degrees is an approximate symmetry. Although there is some variability between runs, Figure 13 highlights we are able to consistently discover rotation coset representatives among the top $24$ cosets.
@@ D.4 Comparison with LieGG in PDE Experiment
For simplicity, we consider the setting where LieGG is given access to the ground truth partial differential equation instead of a predictor network. We also do not perform time stepping and focus solely on the global symmetries of $\frac{\partial u}{\partial t}=\alpha(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}})$. This setup, while slightly different from the experiments for our method, only makes it easier for LieGG to learn the symmetry. The singular values of the discovered generators were shown in Figure 8.
@@
D.5 Coset Discovery of Complex Component Groups
Next, we verify that the coset discovery algorithm of AtlasD can discover more complex component groups. In particular, we search for the symmetries of the function $f(x,y)=|x|+|y|$. The ground truth symmetry group is $D_{4}$, which contains $8$ elements.
We seed the discovery algorithm with $K=256$ representatives and report the unique cosets among the top $q=128$. In Figure 17, we show that AtlasD finds exactly the $8$ elements of $D_{4}$.
D.6 Discovery under Sheared Charts
To further verify the resilience of AtlasD under the choice of atlas, we repeat the PDE experiment under a third atlas where one of the coordinate charts is partially sheared (Figure 18).
We discover a single generator $\begin{bmatrix}-0.368&-1.035\\ 1.101&0.386\end{bmatrix}$ and both cosets. While the noisy chart does worsen the learned generator, it remains recognizable as a rotation.
Figure 17: Discovered symmetry group for $|x|+|y|$
Figure 18: Atlas with sheared chart